# On polynomial interpolation in the monomial basis

### v4

@@ 2 Polynomial interpolation in the monomial basis
In practice, one can take $\{Q_{n}\}_{n=0,1,\dots,N}$ to be a finite sequence of interpolating polynomials $\{P_{n}\}_{n=0,1,\dots,N}$ of $F$ for sets of collocation points with small Lebesgue constants. When the Lebesgue constant $\Lambda_{N}$ is small, it follows from Theorem 2.6 that the condition (50) is satisfied when $\rho_{*}^{N}\lesssim\frac{1}{u}$. We assume here that $N$ is sufficiently small so that $\rho_{*}^{N}\lesssim\frac{1}{u}$. Without loss of generality, we assume that the upper bound for ${\lVert F-P_{n}\rVert}_{L^{\infty}(\Gamma)}$, i.e., $C_{N}\rho_{*}^{-n}$, is tight, in the sense that there exists some integer $n\in[0,N]$ such that ${\lVert F-P_{n}\rVert}_{L^{\infty}(\Gamma)}=C_{N}\rho_{*}^{-n}$. Note that the smallest uniform approximation error we can hope to obtain in practice is $u\cdot{\lVert F\rVert}_{L^{\infty}(\Gamma)}$.
@@
When $u\cdot C_{N}\lesssim\max({\lVert F-P_{N}\rVert}_{L^{\infty}(\Gamma)},u\cdot{\lVert F\rVert}_{L^{\infty}(\Gamma)})$, the use of a monomial basis for interpolation introduces essentially no extra error. Interestingly, this happens both if the polynomial interpolation error decays quickly and if the polynomial interpolation error decays slowly. Suppose that the polynomial interpolation error decays quickly, so that the bound is tight for $n=0$, i.e., ${\lVert F-P_{0}\rVert}_{L^{\infty}(\Gamma)}=C_{N}$. Since $C_{N}\lesssim 2{\lVert F\rVert}_{L^{\infty}(\Gamma)}$, we see that the extra error caused by the use of a monomial basis is bounded by $u\cdot C_{N}\lesssim 2u\cdot{\lVert F\rVert}_{L^{\infty}(\Gamma)}\lesssim u\cdot{\lVert F\rVert}_{L^{\infty}(\Gamma)}$. Examples of this situation are illustrated in Figure 7. Suppose now that the polynomial interpolation error decays slowly, so that bound is tight for $n=N$, i.e., ${\lVert F-P_{N}\rVert}_{L^{\infty}(\Gamma)}=C_{N}\rho_{*}^{-N}$. Since we assumed that $\rho_{*}^{N}\lesssim\frac{1}{u}$, it follows that $u\cdot C_{N}\lesssim{\lVert F-P_{N}\rVert}_{L^{\infty}(\Gamma)}$. Examples of this situation are illustrated in Figure 8.
@@
When $u\cdot C_{N}\gtrsim\max({\lVert F-P_{N}\rVert}_{L^{\infty}(\Gamma)},u\cdot{\lVert F\rVert}_{L^{\infty}(\Gamma)})$, stagnation of convergence can occur. In practice, we observe that the extra error caused by the use of the monomial basis, i.e., $u\cdot{\lVert a^{(N)}\rVert}_{2}$, is close to $u\cdot C_{N}$, so the monomial approximation error ${\lVert F-\widehat{P}_{N}\rVert}_{L^{\infty}(\Gamma)}$ generally stagnates at an error level around $u\cdot C_{N}$. Note that a slow decay in ${\lVert F-P_{n}\rVert}_{L^{\infty}(\Gamma)}$ results in a larger value of $C_{N}$, while a fast decay in ${\lVert F-P_{n}\rVert}_{L^{\infty}(\Gamma)}$ favors a smaller final interpolation error ${\lVert F-P_{N}\rVert}_{L^{\infty}(\Gamma)}$. This means that, for stagnation of convergence to occur, the polynomial interpolation error has to exhibit some combination of slow decay followed by fast decay. Furthermore, note that an upper bound for $C_{N}$ is given by $C_{N}\lesssim\rho_{*}^{N}$ (see Remark 2.5). This means that, the smaller the value of $N$, the smaller the maximum possible value of $C_{N}$, and the more rapid the rate of decay in ${\lVert F-P_{n}\rVert}_{L^{\infty}(\Gamma)}$ required for stagnation of convergence to occur. We present examples of this situation in Figure 9.
(a) $\cos(2x+1)$, $C=3\times 10^{0}$
(b) $\frac{1}{x-\sqrt{2}}$, $C=3\times 10^{0}$
@@
Figure 7: The extra error caused by the use of a monomial basis is negligible. The pink region denotes the bound for $u\cdot{\lVert a^{(n)}\rVert}_{2}$
(a) $\frac{1}{x-0.5i}$, $C=5\times 10^{7}$
(b) ${\lvert x+0.1\rvert}^{2.5}$, $C=5\times 10^{11}$
@@
Figure 8: The extra error caused by the use of a monomial basis is no larger than ${\lVert F-P_{N}\rVert}_{L^{\infty}(\Gamma)}$. The pink region denotes the bound for $u\cdot{\lVert a^{(n)}\rVert}_{2}$
(a) $\cos(8x+1)$, $C=1\times 10^{3}$
@@
Figure 9: Stagnation of convergence. The pink region denotes the bound for $u\cdot{\lVert a^{(n)}\rVert}_{2}$.
@@ 2.2 Practical use of a monomial basis for interpolation
What are the restrictions on polynomial interpolation in the monomial basis? Firstly, extremely high-order global interpolation is impossible in the monomial basis, because the order $N$ must satisfy ${\lVert(V^{(N)})^{-1}\rVert}_{2}\lesssim\frac{1}{u}$ for our estimates to hold, where $u$ denotes machine epsilon. In fact, even if this condition were not required, there would still be no benefit in taking an order larger than this threshold. This is because, in almost all situations, the extra error caused by the use of the monomial basis dominates the monomial approximation error when ${\lVert(V^{(N)})^{-1}\rVert}_{2}>\frac{1}{u}$, leading to a stagnation of convergence.
@@
On the other hand, piecewise polynomial interpolation in the monomial basis over a partition of $\Gamma$ can be carried out stably, provided that the maximum order of approximation over each subpanel is maintained below the threshold $\operatorname*{arg\,max}_{N}{\lVert(V^{(N)})^{-1}\rVert}_{2}\lesssim\frac{1}{u}$, and that the size of $u\cdot{\lVert a^{(N)}\rVert}_{2}\approx u\cdot{\lVert\widehat{a}^{(N)}\rVert}_{2}$ is kept below the size of the polynomial interpolation error, where $a^{(N)}$ and $\widehat{a}^{(N)}$ denote the exact and the computed monomial coefficient vectors, respectively. As demonstrated in Section 2.1, the latter requirement is often satisfied automatically, and when it is not, adding an extra level of subdivision almost always resolves the issue. In addition, the extra error caused by the use of a monomial basis can always be estimated promptly during computation, using the value of $u\cdot{\lVert\widehat{a}^{(N)}\rVert}_{2}$.
@@
Based on the discussion above, we summarize the proper way of using the monomial basis for interpolation as follows. For simplicity, we use the same order of approximation over each subpanel, and denote the order using $N$. Firstly, $N$ needs to be smaller than the threshold $\operatorname*{arg\,max}_{N}{\lVert(V^{(N)})^{-1}\rVert}_{2}\lesssim\frac{1}{u}$. Then, given a function $F:\Gamma\to\mathbbm{C}$ and an error tolerance $\varepsilon$, we subdivide the domain $\Gamma$ until $F$ can be approximated by a polynomial of degree less than $N$ over each subpanel to within an error of $\varepsilon$. Finally, we subdivide the panels further until the norm of the monomial coefficients is less than $\varepsilon/u$ over each subpanel.
@@
Since the convergence rate of piecewise polynomial approximation is $\mathcal{O}(h^{N+1})$, where $h$ and $N$ denote the maximum diameter and minimum order of approximation over all subpanels, respectively, and since the aforementioned threshold is generally not small (e.g., the threshold is approximately equal to $43$ when $\Gamma=[-1,1]$), piecewise polynomial interpolation in the monomial basis converges rapidly so long as we set the value of $N$ to be large enough. Therefore, there is no need to avoid the use of a monomial basis when it offers an advantage over other bases.
@@
It takes $\mathcal{O}(N^{3})$ operations to solve a Vandermonde system of size $N\times N$ by a standard backward stable solver, e.g., LU factorization with partial pivoting. Since the order of approximation $N$ is almost always not large, the solution to the Vandermonde matrix can be computed accurately, in the sense that $\gamma_{N}$ is small, and rapidly, using highly optimized linear algebra libraries, e.g., LAPACK. There also exist specialized algorithms that solve Vandermonde systems in $\mathcal{O}(N^{2})$ operations, e.g., the Björck-Pereyra algorithm [9], the Parker-Traub algorithm [15].
Observation 2.7.
What happens when the order of approximation exceeds the threshold? We observe that, despite that our theory is no longer applicable, the monomial approximation error does not become much larger than the error at the threshold, when the columns of the Vandermonde matrix are ordered as in (20) and when the system is solved by MATLAB’s backslash operator (which implements LU factorization with partial pivoting).
@@ 2.3 Interpolation over an interval
In this section, we consider polynomial interpolation in the monomial basis over an interval $\Gamma=[a,b]\subset\mathbbm{R}$. We suggest the use of the Chebyshev points on the interval $[a,b]$ as the collocation points, because of the following two well-known theorems related to Chebyshev approximation.
The theorem below, originally proved in [12], bounds the growth rate of the Lebesgue constant for the Chebyshev points.
Theorem 2.7.
Let $\Lambda_{N}$ be the Lebesgue constant for the $(N+1)$ Chebyshev points on an interval $[a,b]$. For any nonnegative integer $N$, the Lebesgue constant $\Lambda_{N}$ satisfies $\Lambda_{N}\leq\frac{2}{\pi}\log(N+1)+1$.
The following theorem provides a sufficient condition for the Chebyshev interpolant of a function to converge geometrically. The proof can be found in, for example, Theorem 8.2 in [27]. Recall that the level set $E_{\rho}$ for an interval $[a,b]$ is a Bernstein ellipse with parameter $\rho$, with foci at $a$ and $b$ (see Figure 5(a)).
Theorem 2.8.
Suppose that $F:[a,b]\to\mathbbm{C}$ is analytically continuable to the region $E_{\rho}^{o}$ (see Definition 2.1), and satisfies ${\lVert F\rVert}_{L^{\infty}(E_{\rho}^{o})}\leq M$ for some $M\geq 0$. The $N$th degree Chebyshev interpolant $P_{N}$ of $F$ satisfies
${\lVert F-P_{N}\rVert}_{L^{\infty}([a,b])}\leq\frac{4M}{\rho-1}\rho^{-N},$ (54)
for all $N\geq 0$.
We note that the theorem above is stronger than Theorem 2.3 when $\Gamma$ is an interval, as it specifies the constant factor $C$.
Remark 2.8.
The Legendre points exhibit similar characteristics to the Chebyshev points, and can also be effectively utilized for interpolation over an interval.
@@
In the rest of this section, we provide a series of numerical experiments involving interpolation over intervals. In Figure 10, we report the 2-norm of the inverse of the Vandermonde matrices with Chebyshev collocation points, for the domains $\Gamma=[-1,1]$ and $\Gamma=[0,1]$. Note that when $\Gamma=[-1,1]$, we have that $\rho_{*}=1+\sqrt{2}$ and ${\lVert(V^{(N)})^{-1}\rVert}_{2}\leq\frac{1}{u}$ for $N\leq 43$; when $\Gamma=[0,1]$, we have that $\rho_{*}=3+2\sqrt{2}$ and ${\lVert(V^{(N)})^{-1}\rVert}_{2}\leq\frac{1}{u}$ for $N\leq 22$. In Figure 11, we interpolate functions which can be resolved by a Chebyshev interpolant of degree $N\leq 43$ over $\Gamma=[-1,1]$. In addition to the estimated values of ${\lVert F-P_{N}\rVert}_{L^{\infty}([-1,1])}$ and ${\lVert F-\widehat{P}_{N}\rVert}_{L^{\infty}([-1,1])}$, we plot three additional curves in each figure: the estimated values of $u\cdot{\lVert a^{(N)}\rVert}_{2}$ based on inequality (31), the upper bound $C\rho_{*}^{N}$ for ${\lVert F-P_{N}\rVert}_{L^{\infty}([-1,1])}$, and the upper bound $u\cdot C$ for $u\cdot{\lVert a^{(N)}\rVert}_{2}$. In Figure 12, we provide similar experiments for the case where $\Gamma=[0,1]$. Based on these experimental results, one can observe that the convergence stagnates after the monomial approximation error ${\lVert F-\widehat{P}_{N}\rVert}_{L^{\infty}(\Gamma)}$ reaches $u\cdot{\lVert a^{(N)}\rVert}_{2}$, which implies that inequality (51) is sharp. In addition, the values of $u\cdot{\lVert a^{(N)}\rVert}_{2}$ are always within the upper bound $u\cdot C$, which is inline with our analysis in Section 2.1.
(a) $\Gamma=[-1,1]$
(b) $\Gamma=[0,1]$
@@
Figure 10: The 2-norm of the inverse of a Vandermonde matrix with Chebyshev collocation points over an interval $\Gamma$, and its upper bound, for different orders of approximation. We note that $\rho_{*}=1+\sqrt{2}$ when $\Gamma=[-1,1]$, and $\rho_{*}=3+2\sqrt{2}$ when $\Gamma=[0,1]$.
@@
Figure 11: Polynomial interpolation in the monomial basis over $\Gamma=[-1,1]$. The constant $\rho_{*}$ equals $1+\sqrt{2}$.

### v5

@@ 2 Polynomial interpolation in the monomial basis
In practice, one can take $\{Q_{n}\}_{n=0,1,\dots,N}$ to be a finite sequence of interpolating polynomials $\{P_{n}\}_{n=0,1,\dots,N}$ of $F$ for sets of interpolation points with Lebesgue constants $\lesssim 1$. When the Lebesgue constant $\Lambda_{N}$ is small, it follows from Theorem 2.5 that the condition (34) is satisfied when $\rho_{*}^{N}\lesssim\frac{1}{u}$. Suppose, then, that $N$ is sufficiently small so that $\rho_{*}^{N}\lesssim\frac{1}{u}$. Without loss of generality, we assume that the upper bound for ${\lVert F-P_{n}\rVert}_{L^{\infty}(\Omega)}$, i.e., $C_{N}\rho_{*}^{-n}$, is tight, in the sense that there exists some integer $n\in[0,N]$ such that ${\lVert F-P_{n}\rVert}_{L^{\infty}(\Omega)}=C_{N}\rho_{*}^{-n}$. Note that the smallest uniform approximation error we can hope to obtain in practice is $u\cdot{\lVert F\rVert}_{L^{\infty}(\Omega)}$.
@@
When $u\cdot C_{N}\lesssim\max({\lVert F-P_{N}\rVert}_{L^{\infty}(\Omega)},u\cdot{\lVert F\rVert}_{L^{\infty}(\Omega)})$, the use of a monomial basis for interpolation introduces essentially no extra error. Interestingly, this happens both if the polynomial interpolation error decays quickly and if the polynomial interpolation error decays slowly. Suppose that the polynomial interpolation error decays quickly, so that the bound is tight for $n=0$, i.e., ${\lVert F-P_{0}\rVert}_{L^{\infty}(\Omega)}=C_{N}$. Since $C_{N}\lesssim 2{\lVert F\rVert}_{L^{\infty}(\Omega)}$, we see that the extra error caused by the use of a monomial basis is bounded by $u\cdot C_{N}\lesssim 2u\cdot{\lVert F\rVert}_{L^{\infty}(\Omega)}\lesssim u\cdot{\lVert F\rVert}_{L^{\infty}(\Omega)}$. Examples of this situation are illustrated in Figure 7. Alternatively, suppose that the polynomial interpolation error decays slowly, so that bound is tight for $n=N$, i.e., ${\lVert F-P_{N}\rVert}_{L^{\infty}(\Omega)}=C_{N}\rho_{*}^{-N}$. Since we assumed that $\rho_{*}^{N}\lesssim\frac{1}{u}$, it follows that $u\cdot C_{N}\lesssim{\lVert F-P_{N}\rVert}_{L^{\infty}(\Omega)}$. Examples of this situation are illustrated in Figure 8.
@@
When $u\cdot C_{N}\gtrsim\max({\lVert F-P_{N}\rVert}_{L^{\infty}(\Omega)},u\cdot{\lVert F\rVert}_{L^{\infty}(\Omega)})$, stagnation of convergence can occur. In practice, we observe that the extra error caused by the use of the monomial basis, i.e., $u\cdot{\lVert a^{(N)}\rVert}_{2}$, is close to $u\cdot C_{N}$, so the monomial approximation error ${\lVert F-\widehat{P}_{N}\rVert}_{L^{\infty}(\Omega)}$ generally stagnates at an error level around $u\cdot C_{N}$. Note that a slow decay in ${\lVert F-P_{n}\rVert}_{L^{\infty}(\Omega)}$ results in a larger value of $C_{N}$, while a fast decay in ${\lVert F-P_{n}\rVert}_{L^{\infty}(\Omega)}$ favors a smaller final interpolation error ${\lVert F-P_{N}\rVert}_{L^{\infty}(\Omega)}$. This means that, for stagnation of convergence to occur, the polynomial interpolation error has to exhibit some combination of slow decay followed by fast decay. Furthermore, note that an upper bound for $C_{N}$ is given by $C_{N}\lesssim\rho_{*}^{N}$ (see Remark 2.6). This means that, the smaller the value of $N$, the smaller the maximum possible value of $C_{N}$, and the more rapid the rate of decay in ${\lVert F-P_{n}\rVert}_{L^{\infty}(\Omega)}$ required for stagnation of convergence to occur. We present examples of this situation in Figure 9.
(a) $F(x)=\cos(2x+1)$, $C=3\times 10^{0}$
(b) $F(x)=\frac{1}{x-\sqrt{2}}$, $C=3\times 10^{0}$
@@
Figure 7: Cases when the extra error caused by the use of a monomial basis is negligible. The function $F$ is interpolated over $\Omega=[-1,1]$. The value of $C$ is chosen such that ${\lVert F-P_{N}\rVert}_{L^{\infty}([-1,1])}\leq C\rho_{*}^{-N}$ for $N=0,1,...,40$. We highlight the region bounded above by $u\cdot C$ in pink.
(a) $F(x)=\frac{1}{x-0.5i}$, $C=5\times 10^{7}$
(b) $F(x)={\lvert x+0.1\rvert}^{2.5}$, $C=5\times 10^{11}$
@@
Figure 8: Cases when the extra error caused by the use of a monomial basis is no larger than ${\lVert F-P_{N}\rVert}_{L^{\infty}(\Omega)}$. The function $F$ is interpolated over $\Omega=[-1,1]$. The value of $C$ is chosen such that ${\lVert F-P_{N}\rVert}_{L^{\infty}([-1,1])}\leq C\rho_{*}^{-N}$ for $N=0,1,...,40$. We highlight the region bounded above by $u\cdot C$ in pink.
(a) $F(x)=\cos(8x+1)$, $C=1\times 10^{3}$
@@
Figure 9: Cases when convergence stagnates. The function $F$ is interpolated over $\Omega=[-1,1]$. The value of $C$ is chosen such that ${\lVert F-P_{N}\rVert}_{L^{\infty}([-1,1])}\leq C\rho_{*}^{-N}$ for $N=0,1,...,40$. We highlight the region bounded above by $u\cdot C$ in pink.
@@ 2.2 Practical use of a monomial basis for interpolation
What are the restrictions on polynomial interpolation in the monomial basis? Firstly, extremely high-order global interpolation is impossible in the monomial basis, because the order $N$ must satisfy ${\lVert(V^{(N)})^{-1}\rVert}_{2}\lesssim\frac{1}{u}$ for our estimates to hold, where $u$ denotes machine epsilon. In fact, even if this condition were not required, there would still be no benefit in taking an order larger than this threshold. For most functions, if a polynomial of degree exceeding the threshold is needed to resolve the function to machine precision, the error caused by the use of a monomial basis typically dominates the true polynomial interpolation error, which leads to stagnation of convergence. This can be seen from the discussion in Section 2.1.
@@
On the other hand, piecewise polynomial interpolation in the monomial basis over a partition of $\Omega$ can be carried out stably, provided that the maximum order of approximation over each subelement of $\Omega$ is maintained below the threshold $\operatorname*{arg\,max}_{N}{\lVert(V^{(N)})^{-1}\rVert}_{2}\lesssim\frac{1}{u}$, and that the size of $u\cdot{\lVert a^{(N)}\rVert}_{2}\approx u\cdot{\lVert\widehat{a}^{(N)}\rVert}_{2}$ is kept below the size of the polynomial interpolation error, where $a^{(N)}$ and $\widehat{a}^{(N)}$ denote the exact and the computed monomial coefficient vectors, respectively. As demonstrated in Section 2.1, the latter requirement is often satisfied automatically, and when it is not, adding an extra level of subdivision almost always resolves the issue. In addition, the extra error caused by the use of a monomial basis can always be estimated promptly during computation, using the value of $u\cdot{\lVert\widehat{a}^{(N)}\rVert}_{2}$.
@@
Based on the discussion above, we summarize a proper method of using the monomial basis for interpolation, as follows. For simplicity, we use the same order of approximation over each subelement, and denote the order by $N$. This value of $N$ needs to be smaller than the threshold $\operatorname*{arg\,max}_{N}{\lVert(V^{(N)})^{-1}\rVert}_{2}\lesssim\frac{1}{u}$. Given a function $F:\Omega\to\mathbbm{C}$ and an error tolerance $\varepsilon$, we subdivide the domain $\Omega$ until $F$ can be approximated by a polynomial of degree less than or equal to $N$ over each subelement to within an error of $\varepsilon$. Finally, we subdivide the subelements further until the 2-norm of the monomial coefficient vector $\widehat{a}^{(N)}$ is less than $\varepsilon/u$ over each subelement.
@@
Since the convergence rate of piecewise polynomial approximation is $\mathcal{O}(h^{N+1})$, where $h$ and $N$ denote the maximum diameter and minimum order of approximation over all subelements, respectively, and since the aforementioned threshold is generally not small (e.g., the threshold is approximately equal to $43$ when $\Omega=[-1,1]$), piecewise polynomial interpolation in the monomial basis converges rapidly so long as we set the value of $N$ to be large enough. Therefore, there is no need to avoid the use of a monomial basis when it offers an advantage over other bases.
@@
It takes $\mathcal{O}(N^{3})$ operations to solve a Vandermonde system of size $N\times N$ by a standard backward stable solver. Since the order of approximation $N$ is almost always not large, the solution to the Vandermonde system can be computed both rapidly, as well as accurately in the sense that $\gamma_{N}$ is small, using highly optimized linear algebra libraries, e.g., LAPACK. There also exist specialized algorithms that solve Vandermonde systems in $\mathcal{O}(N^{2})$ operations, e.g., the Björck-Pereyra algorithm [10] and the Parker-Traub algorithm [16]. We note that the Parker-Traub algorithm is known to be backward stable [16], while the stability of the Björck-Pereyra algorithm is known only under very restrictive conditions [20].
3 Numerical experiments
@@ 3.1 Interpolation over an interval
In this section, we consider polynomial interpolation in the monomial basis over an interval $\Omega=[a,b]\subset\mathbbm{R}$. We compute the polynomial interpolants using Chebyshev points on the interval $[a,b]$, since their associated Lebesgue constant, $\Lambda_{N}$, is bounded by $\frac{2}{\pi}\log(N+1)+1$, where $N$ is the order of approximation [12].
In Figure 10, we report the values of ${\lVert(V^{(N)})^{-1}\rVert}_{2}$ and its upper bound $\rho_{*}^{N}\Lambda_{N}$ (see Theorem 2.5), for the domains $\Omega=[-1,1]$ and $\Omega=[0,1]$. One can see that the upper bound is close to being tight. Note that when $\Omega=[-1,1]$, we have that $\rho_{*}=1+\sqrt{2}$ and ${\lVert(V^{(N)})^{-1}\rVert}_{2}\leq\frac{1}{u}$ for $N\leq 43$; when $\Omega=[0,1]$, we have that $\rho_{*}=3+2\sqrt{2}$ and ${\lVert(V^{(N)})^{-1}\rVert}_{2}\leq\frac{1}{u}$ for $N\leq 22$.
@@
In Figure 11, we interpolate several functions over $\Omega=[-1,1]$. We denote the interpolant constructed using the monomial basis by $\widehat{P}_{N}$, and estimate the true polynomial interpolant $P_{N}$ using the Lagrange basis evaluated by the Barycentric interpolation formula. In addition to the estimated values of ${\lVert F-P_{N}\rVert}_{L^{\infty}([-1,1])}$ and ${\lVert F-\widehat{P}_{N}\rVert}_{L^{\infty}([-1,1])}$, we plot three additional curves in each figure: the estimated values of $u\cdot{\lVert a^{(N)}\rVert}_{2}$ based on inequality (15), the upper bound $C\rho_{*}^{-N}$ for ${\lVert F-P_{N}\rVert}_{L^{\infty}([-1,1])}$, where $C$ was chosen so that ${\lVert F-P_{N}\rVert}_{L^{\infty}([-1,1])}\leq C\rho_{*}^{-N}$ for $N=0,1,...,40$, and the approximate upper bound $u\cdot C$ for $u\cdot{\lVert a^{(N)}\rVert}_{2}$. In Figure 12, we provide similar experiments for the case where $\Omega=[0,1]$, this time choosing $C$ so that ${\lVert F-P_{N}\rVert}_{L^{\infty}([0,1])}\leq C\rho_{*}^{-N}$ for $N=0,1,...,20$. Based on these experimental results, one can observe that the convergence stagnates after the monomial approximation error ${\lVert F-\widehat{P}_{N}\rVert}_{L^{\infty}(\Omega)}$ reaches $u\cdot{\lVert a^{(N)}\rVert}_{2}$, which implies that inequality (35) is sharp. Moreover, the values of $u\cdot{\lVert a^{(N)}\rVert}_{2}$ are always less than the upper bound $u\cdot C$, which agrees with our analysis in Section 2.1.
What happens when the order of approximation $N$ exceeds the threshold, i.e., when ${\lVert(V^{(N)})^{-1}\rVert}_{2}\geq\frac{1}{u}$? In Figure 13, we use two different solvers to solve the Vandermonde system with Chebyshev interpolation points over the interval $[-1,1]$: LU factorization with partial pivoting, and the truncated singular value decomposition (TSVD) with a truncation cutoff of $10^{-14}$. Since backward stability no longer guarantees that ${\lVert\widehat{a}^{(N)}\rVert}_{2}\approx{\lVert a^{(N)}\rVert}_{2}$ when $N$ is greater than the threshold, the upper bound for the monomial approximation error (16) becomes invalid, and our previous analysis is no longer applicable. As shown in the experiments, the approximation error does not improve when $N>43$. When the system is solved using LU decomposition, the error fluctuates. We also note that the error could grow exponentially after $N$ exceeds the threshold, depending on the particular choice of the backward stable solver. When the TSVD is used, the error is stable. This can be explained by Theorem 2.1 in [37], which provides an analogous bound to (16) for the case of the TSVD.
(a) $\Omega=[-1,1]$
(b) $\Omega=[0,1]$
@@
Figure 10: The 2-norm of $(V^{(N)})^{-1}$ with Chebyshev interpolation points over an interval $\Omega$, and its upper bound, for different orders of approximation. We note that $\rho_{*}=1+\sqrt{2}$ when $\Omega=[-1,1]$, and $\rho_{*}=3+2\sqrt{2}$ when $\Omega=[0,1]$.
@@
Figure 11: Polynomial interpolation in the monomial basis over $\Omega=[-1,1]$. The constant $\rho_{*}$ equals $1+\sqrt{2}$. The value of $C$ is chosen such that ${\lVert F-P_{N}\rVert}_{L^{\infty}([-1,1])}\leq C\rho_{*}^{-N}$ for $N=0,1,...,40$. We highlight the region bounded above by $u\cdot C$ in pink.